Meeting:2011/07/07

From SOWNWiki
Jump to: navigation, search

Meeting (to be) held on 2011/07/07 at 18:30 in Access Grid Room

Previous meeting: 30 June 2011 18:30:00 Next meeting: 14 July 2011 19:00:00


Minutes

Amateur Radio (18:30-19:00)

  • Rallies (mike)
    • IVARC 8/7/11


  • Bunfight Ideas
    • "What is Amateur Radio" leaflet


  • AOB


Joint Agenda (19:00-19:15)


  • Bank account status (Savetaz)
    • Membership fees collection
    • RSGB affiliation



SOWN (19:15-20:00)

  • Admin Tasks
    • VoIP card requirements (SjH)
    • New Auth SSL certificate roll-out (Davetaz)

Auth SSL certificate rolled out, no major problems. RADIUS certificate needs updating to be the same as the Website, we have about 2 months before the current cert expires.

    • Move SOWN and trunk ports to the Access Grid room (daveruss)

No progress

No progress and SjH/morse not present. Mike has been having fun with Extreme routers and their virtual routers.

Request has been accepted for the new semester.

    • Chamberlain

daveruss has been in contact with Daz and he has not had time to look at this as of yet.


All surveys complete and laptops are being returned. NOC results have been made available to Mr Norman of the NOC IT department. Mike has confirmed his position as a slave of iSolutions. The project has been a positive demonstration of how iSolutions can work with ECS and research interests to produce a positive result. It is hoped that this can be continued on as we work towards a campus wide Eduroam deployment.

This is now complete and is working. Positive comments have been made with regards to its ability to quickly portray the current health of the ECS/SOTON/SOWN radius deployments. Currently it looks good for iSolutions to switch from their current RADIUS servers to their new RADIUS servers. ECS unfortunately is not is such a healthy state.

Leth not present

    • Rebuilding VPN

On hold

Nothing to say

Nothing to day


  • AOB

Janet node came up today. It was found out that we were only syslogging the Kernel log. We are now logging more logs via syslogng. All nodes that were up also have this update. Nodes are logging back to VPN.

Janet node had some network issues as a result of some changes performed on it however it is now operating again although not quite as Janet would like it to be. To prevent this in the future there is now the ability to prevent certain scripts being run on on particular nodes.

Facts about "2011/07/07"
Has date7 July 2011 18:30:00 +
Has end date7 July 2011 20:00:00 +
Has location32/3073 +